OpenStudy (natasha18):

When graphing parabolas I know there's a specific way to tell whether to shade the inside or the outside. It depends on which sign you use in your equation(<,>) One means to shade the inside and one means to shade the outside. Does anyone know which goes with which??

OpenStudy (hossamhoussien):

I think this might helps http://www.purplemath.com/modules/grphquad2.htm

OpenStudy (natasha18):

Thanks! :)

OpenStudy (danjs):

greater than , take any point and see which side of the thing the larger value is, shade above for greater than , and below for less than

OpenStudy (mathmale):

If all you're doing is graphing a parabola, you don't shade any part of the graph. If you were told to shade, then you are dealing with an inequality problem.

OpenStudy (danjs):

inside would be the greater than part it opens wider forever, you can say ' above' and 'below' for that , i think

OpenStudy (danjs):

inside the parabola is always a larger value than the value on the curve, for y=x^2

OpenStudy (danjs):

if it is y = -x^2, it opens down, things are opposite, inside that one will be the less than smaller values

OpenStudy (mathmale):

Again, where you shade a parabola depends upon the particular inequality you're graphing. Where would you shade\[5x^2-2x +1\le8?\]

OpenStudy (mathmale):

Once again: NO SHADING unless you are dealing with an inequality.

OpenStudy (danjs):

Open Up +x^2 greater than values are inside the curve, above less than is everything else, below opens down -x^2 opposite case greater than values are outside the curve, above less than values are inside the curve, below

OpenStudy (mathmale):

the parabola represented by \[5x^2-2x +1\le8\] opens up. We know that much. Note that this is an inequality. Which line or curve represents the smallest possible y values? the largest possible y valule?

OpenStudy (danjs):

for any function if it is greater than , the solution region is above the curve if it is less than , then the solution region is below the curve

OpenStudy (natasha18):

Thanks for all the help guys! I understand it much better now :)

OpenStudy (mathmale):

Dan: Please take a look at this product of Wolfram Alpha: http://www.wolframalpha.com/input/?i=5x%5E2-2x+%2B+1+is+less+than+or+equal+to+8

OpenStudy (danjs):

that is not a function y=f(x)

OpenStudy (danjs):

solve the thing for y if y > shade above the curve y < shade below

OpenStudy (danjs):

oh, that is the solution set for the parabola and the line

OpenStudy (danjs):

y > parabola and y < line results in that region, and in that case you can combine the two things to a single inequality parabola < y < line that is how you write that sort of inequality parabola < line

OpenStudy (danjs):

For just one function , all the area above is greater , all area below is less than compound functions like yours, has shading only where both the parts solution regions overlap and are in common